OpenStudy (anonymous):

Solve for n. 8/3/4 + n + 2/3/8 + 4/1/2 = 23 and 1/8

OpenStudy (anonymous):

Wanna make sure I did it right!

OpenStudy (anonymous):

no

OpenStudy (anonymous):

lol no thats apart of the problem = 23/1/8

OpenStudy (anonymous):

If you are having a hard time understanding how to do the problems, here is a step by step look at how I did it. \[8\frac{ 3 }{ 4 } + n + 2\frac{ 3 }{ 8 } + 4\frac{ 1 }{ 2 } = 23\frac{ 1 }{ 8 }\] That's the problem, right? If so, you need to take all the whole numbers and turn them into their fraction equivalents. The problem would look like: \[\frac{ 35 }{ 4 } + n + \frac{ 19 }{ 8 } + \frac{ 9 }{ 2 } = \frac{ 185 }{ 8 }\] You need to simplify all the denominators (bottom numbers). The number 8 is the lowers common denominator we can use. You will need to multiply the fractions by 2's and 4's. The problem then would look like: \[\frac{ 70 }{ 8 } + n + \frac{ 19 }{ 8 } + \frac{ 36 }{ 8 } = \frac{ 185 }{ 8 }\] Now you can add them, because they have similar denominators. You would get: \[n + \frac{ 125 }{ 8 } = \frac{ 185 }{ 8 }\] Then just subtract and simplify! \[n = \frac{ 185 }{ 8 }\ - \frac{ 125 }{ 8 }\] \[n = \frac{ 60 }{ 8 } = 7\frac{ 4 }{ 8 } = 7\frac{ 1 }{ 2 }\] I think this is right, I hope I helped clear things up.

OpenStudy (anonymous):

I might be wrong, I don't know, finals have fried my brain.